# Harnessing the Universal Geometry of Embeddings

### v2

@@ 7 Related work
Representation alignment. Similarities between representations of different neural networks are investigated in [24, 28, 52, 4, 17]. CCA [36], SVCCA, [45], CKA [21], ICA [55], and GUI-based systems such as embComp [15] compare embeddings from different subspaces. Methods like [33, 39, 34, 42] harness representation similarity for zero-shot stitching, substitution, and multi-modal adaptation. All rely on some amount of paired data, which is difficult to reduce [5]. Our method does not just measure similarity, we learn how to translate representations across spaces without any paired data.
Unsupervised transport. The problem of unsupervised transport has been studied for images [16, 32, 61], words [54, 9, 14, 8], and natural language sequences [46, 25, 1, 3, 56]. Our method builds on these works, which often employ a combination of cycle-consistency and adversarial loss. [48] proposes a solver for matching small sets of embeddings between different vision-language models. Our method goes well beyond matching by taking unknown embeddings and generating matching embeddings in the space of another model.
@@ Appendix A Compute
  • Optimal assignment baselines (36 models) were run exclusively on CPU. We implemented 4 different assignment algorithms, each requiring approximately 4 hours of computation time. The Gromov-Wasserstein algorithm was particularly computationally intensive and was included for CLIP due to dimensional differences.
In total, our experiments consumed approximately 180 GPU days for training and an additional 78 GPU hours for evaluation and analysis. There were approximately 144 CPU hours required for optimal assignment.
Appendix B Full out-of-distribution translation results
@@
vec2vec Naïve Optimal Transport
$E_{1}$ $E_{2}$ $\cos(\cdot)$ Top-1 Rank $\cos(\cdot)$ Top-1 Rank Top-1 Rank
gte gtr 0.71 (0.0) 0.95 1.29 (0.1) 0.04 (0.0) 0.00 386.58 (8.3) 0.00 383.41 (8.1)
stel. 0.86 (0.0) 1.00 1.00 (0.0) 0.58 (0.0) 1.00 1.00 (0.0) 1.00 1.00 (0.0)
e5 0.83 (0.0) 0.91 1.57 (0.2) 0.68 (0.0) 1.00 1.00 (0.0) 1.00 1.00 (0.0)
gran. 0.73 (0.0) 0.94 1.33 (0.1) 0.00 (0.0) 0.00 408.81 (8.3) 0.01 398.16 (8.2)
gtr gte 0.85 (0.0) 0.96 1.29 (0.2) 0.04 (0.0) 0.00 392.01 (8.2) 0.00 384.00 (8.1)
stel. 0.77 (0.0) 0.96 1.10 (0.0) 0.00 (0.0) 0.00 394.69 (8.3) 0.00 390.67 (8.2)
e5 0.80 (0.0) 0.53 13.38 (1.2) 0.03 (0.0) 0.00 400.85 (8.2) 0.00 400.48 (8.2)
gran. 0.79 (0.0) 0.98 2.41 (0.6) -0.04 (0.0) 0.00 411.53 (8.3) 0.00 401.01 (8.2)
stel. gte 0.90 (0.0) 1.00 1.00 (0.0) 0.58 (0.0) 1.00 1.00 (0.0) 1.00 1.00 (0.0)
gtr 0.77 (0.0) 1.00 1.00 (0.0) 0.00 (0.0) 0.00 393.07 (8.1) 0.00 390.26 (8.2)
e5 0.85 (0.0) 0.98 1.05 (0.0) 0.37 (0.0) 0.89 1.55 (0.1) 1.00 1.00 (0.0)
gran. 0.79 (0.0) 0.99 1.09 (0.1) 0.00 (0.0) 0.00 418.16 (8.4) 0.00 400.95 (8.2)
e5 gte 0.87 (0.0) 0.99 1.02 (0.0) 0.68 (0.0) 1.00 1.00 (0.0) 1.00 1.00 (0.0)
gtr 0.67 (0.0) 0.80 3.10 (0.6) 0.03 (0.0) 0.00 401.16 (8.4) 0.00 400.49 (8.2)
stel. 0.75 (0.0) 0.98 1.06 (0.0) 0.37 (0.0) 1.00 1.00 (0.0) 1.00 1.00 (0.0)
gran. 0.79 (0.0) 0.98 1.08 (0.0) 0.02 (0.0) 0.00 405.75 (8.3) 0.00 401.00 (8.2)
gran. gte 0.85 (0.0) 0.95 1.26 (0.1) 0.00 (0.0) 0.00 406.73 (8.2) 0.01 398.14 (8.2)
gtr 0.74 (0.0) 0.99 1.09 (0.1) -0.04 (0.0) 0.00 415.61 (8.2) 0.00 400.99 (8.2)
stel. 0.77 (0.0) 0.96 1.11 (0.0) 0.00 (0.0) 0.00 417.27 (8.2) 0.00 400.97 (8.2)
e5 0.83 (0.0) 0.87 3.10 (0.7) 0.02 (0.0) 0.00 405.53 (8.1) 0.00 401.00 (8.2)
@@
Table 6: Out-of-distribution translations on TweetTopic (with baselines): vec2vec models trained on NQ and evaluated on the entire TweetTopic test set (800 tweets). The rank metric varies from 1 to 800, thus 400 corresponds to a random ordering. Since Optimal Assignment is a matching method, cosine distances are not applicable. Standard errors are shown in parentheses.
vec2vec Naïve Optimal Transport
$E_{1}$ $E_{2}$ $\cos(\cdot)$ Top-1 Rank $\cos(\cdot)$ Top-1 Rank Top-1 Rank
gte gtr 0.69 (0.0) 0.12 256.63 (6.4) 0.08 (0.0) 0.00 4229.90 (26.2) 0.00 4096.63 (26.1)
stel. 0.85 (0.0) 1.00 1.00 (0.0) 0.56 (0.0) 1.00 1.00 (0.0) 1.00 1.00 (0.0)
e5 0.86 (0.0) 0.54 17.71 (0.9) 0.69 (0.0) 0.98 1.04 (0.0) 1.00 1.00 (0.0)
gran. 0.73 (0.0) 0.09 342.15 (7.8) 0.01 (0.0) 0.00 3946.19 (25.8) 0.00 3802.92 (25.9)
gtr gte 0.84 (0.0) 0.12 279.56 (6.9) 0.08 (0.0) 0.00 4180.47 (26.2) 0.00 4096.99 (26.1)
stel. 0.72 (0.0) 0.27 127.92 (4.4) 0.00 (0.0) 0.00 4296.04 (26.1) 0.00 4096.53 (26.1)
e5 0.82 (0.0) 0.01 1413.80 (18.3) 0.09 (0.0) 0.00 4064.47 (26.2) 0.00 4010.13 (26.1)
gran. 0.78 (0.0) 0.51 35.27 (1.9) -0.02 (0.0) 0.00 4023.67 (26.1) 0.00 3964.83 (26.1)
stel. gte 0.91 (0.0) 1.00 1.00 (0.0) 0.56 (0.0) 1.00 1.00 (0.0) 1.00 1.00 (0.0)
gtr 0.75 (0.0) 0.56 17.70 (1.0) 0.00 (0.0) 0.00 4339.83 (26.2) 0.00 4097.00 (26.1)
e5 0.85 (0.0) 0.51 26.33 (1.2) 0.35 (0.0) 0.59 12.68 (0.6) 1.00 1.00 (0.0)
gran. 0.77 (0.0) 0.14 221.95 (5.9) -0.01 (0.0) 0.00 3951.42 (25.9) 0.00 3776.52 (26.0)
e5 gte 0.87 (0.0) 0.60 32.59 (2.6) 0.69 (0.0) 0.98 1.09 (0.0) 1.00 1.00 (0.0)
gtr 0.66 (0.0) 0.01 1029.64 (14.9) 0.09 (0.0) 0.00 4032.85 (26.2) 0.00 4010.06 (26.1)
stel. 0.75 (0.0) 0.46 32.12 (1.4) 0.35 (0.0) 0.86 2.49 (0.1) 1.00 1.01 (0.0)
gran. 0.78 (0.0) 0.21 151.09 (4.6) 0.02 (0.0) 0.00 4008.10 (25.9) 0.00 3932.58 (26.2)
gran. gte 0.85 (0.0) 0.08 346.21 (7.8) 0.01 (0.0) 0.00 3978.35 (26.1) 0.00 3808.18 (25.9)
gtr 0.74 (0.0) 0.60 23.38 (1.6) -0.02 (0.0) 0.00 4010.00 (25.8) 0.00 3962.83 (26.1)
stel. 0.72 (0.0) 0.13 242.23 (6.1) -0.01 (0.0) 0.00 3900.74 (26.2) 0.00 3780.44 (26.0)
e5 0.84 (0.0) 0.12 361.06 (8.7) 0.02 (0.0) 0.00 4024.92 (26.1) 0.00 3937.63 (26.2)
@@
Table 7: Out-of-distribution translations on MIMIC (with baselines): vec2vec models trained on NQ and evaluated on an 8192-record subset of MIMIC. The rank metric varies from 1 to 8192, thus 4096 corresponds to a random ordering. Since Optimal Assignment is a matching method, cosine distances are not applicable. Standard errors are shown in parentheses.
Appendix C Ablations
C.1 Ablating components of our method
@@
Table 8: gte $\to$ gtr translators trained without individual components of our method on NQ and evaluated on a 65536-text subset of NQ (chunked in batches of 8192). The rank metric varies from 1 to 8192, thus 4096 corresponds to a random ordering. Since Optimal Assignment is a matching method, cosine distances are not applicable. Standard errors are shown in parentheses.
C.2 Amount of data needed to learn translation
@@
Table 9: gte $\to$ gtr translators trained with different amounts of gte data: vec2vec models trained on NQ and evaluated an 8192-record subset of NQ. The rank metric varies from 1 to 8192, thus 4096 corresponds to a random ordering. Since Optimal Assignment is a matching method, cosine distances are not applicable. Standard errors are shown in parentheses.
@@
Appendix D Prompt for measuring information extraction

### v3

@@ 7 Related work
Representation alignment. Similarities between representations of different neural networks are investigated in [26, 31, 58, 5, 18, 60, 30]. Methods based on CCA [41], SVCCA, [50], CKA [23, 37], ICA [62], time-series [38], and GUIs [16] have been used to compare embeddings from different subspaces. [36, 44, 39, 56, 47] harness representation similarity for zero-shot stitching, substitution, domain transfer, and multi-modal adaptation. All rely on some amount of paired data, which is difficult to reduce [6]. Our method does not just measure similarity, we learn how to translate representations across spaces without any paired data.
Optimal transport. The problem of unsupervised optimal transport has been studied for image style transfer [17, 35, 68], word translation [61, 10, 15, 9, 20], and natural language sequence translation [51, 27, 1, 4, 63, 3]. Our method builds on these works, which often employ a combination of cycle-consistency and adversarial loss. Importantly, unlike prior word and sequence translation methods, multiple representations of the same input (e.g., heavily overlapping word vocabularies) are unavailable in our setting. [53] proposes a solver for matching small sets of embeddings between different vision-language models. Our method goes well beyond matching by taking unknown embeddings and generating matching embeddings in the space of another model.
@@ Appendix A Compute
  • Optimal transport baselines (36 models and 4 different algorithms) were run exclusively on CPU, each requiring approximately 4 hours of computation time.
In total, our experiments consumed approximately 180 GPU days for training and an additional 78 GPU hours for evaluation and analysis. There were approximately 144 CPU hours required for optimal transport.
Appendix B Extended oracle-aided optimal transport baseline results
Let $u_{i}=M_{1}(d_{i})$ and $v_{i}=M_{2}(d_{i})$ denote embeddings of the same document $d_{i}$ from two different embedding models. In Section˜5, we solve the optimal assignment problem:
$\pi^{*}=\arg\min_{\pi}\sum_{i=1}^{n}\cos(u_{i},v_{\pi(i)}),$
using four algorithms: Hungarian (linear sum assignment), Earth Mover’s Distance (EMD), Sinkhorn, and Gromov-Wasserstein. Note that this baseline computes matchings and transports between embeddings derived from the same underlying texts, strongly favoring optimal transport (OT) methods. Nevertheless, OT still struggles when embeddings originate from different model backbones.
Since the Hungarian algorithm produces a discrete matching, it is evaluated only using Top-1 Accuracy, while the other algorithms are evaluated across all metrics. For each experiment, the lowest-rank solver is reported in Table˜2 and Table˜4 (denoted by symbols in the final column). Evaluation metrics are defined as follows:
  1. 1.
    Top-1 Accuracy: Fraction of embeddings correctly identified as closest pairs, calculated by either selecting the maximum transported mass per embedding or applying the Hungarian algorithm directly to the transport plan $P$. We report the higher accuracy between the two.
  2. 2.
    Mean Rank: Average rank position of the correct embedding match $v_{i}$ when sorted by descending transported mass $P_{ij}$ from $u_{i}$:
    $\text{rank}(v_{i})=\text{position of }v_{i}\text{ among sorted }P_{ij}.$
  3. 3.
    Mean Cosine Similarity: Average cosine similarity between barycenters and true counterparts:
    $v^{\prime}_{i}=\frac{\sum_{j=1}^{n}P_{ij}v_{j}}{\sum_{j=1}^{n}P_{ij}},\quad\text{Similarity}=\frac{1}{n}\sum_{i=1}^{n}\cos(v^{\prime}_{i},v_{i}).$
Appendix C Full out-of-distribution translation results
@@
vec2vec Naïve Baseline OT Baseline
$E_{1}$ $E_{2}$ $\cos(\cdot)$ T-1 Rank $\cos(\cdot)$ T-1 Rank $\cos(\cdot)$ T-1 Rank
gra. gtr 0.74 (0.0) 0.99 1.09 (0.1) -0.04 (0.0) 0.00 415.61 (8.2) 0.51 (0.0) 0.01 398.32 (8.2)
gte 0.85 (0.0) 0.95 1.26 (0.1) 0.00 (0.0) 0.00 406.73 (8.2) 0.74 (0.0) 0.01 398.14 (8.2)
ste. 0.77 (0.0) 0.96 1.11 (0.0) 0.00 (0.0) 0.00 417.27 (8.2) 0.56 (0.0) 0.00 399.57 (8.2)
e5 0.83 (0.0) 0.87 3.10 (0.7) 0.02 (0.0) 0.00 405.53 (8.1) 0.75 (0.0) 0.01 398.33 (8.2)
gtr gra. 0.79 (0.0) 0.98 2.41 (0.6) -0.04 (0.0) 0.00 411.53 (8.3) 0.57 (0.0) 0.01 398.29 (8.2)
gte 0.85 (0.0) 0.96 1.29 (0.2) 0.04 (0.0) 0.00 392.01 (8.2) 0.86 (0.0) 0.00 384.00 (8.1)
ste. 0.77 (0.0) 0.96 1.10 (0.0) 0.00 (0.0) 0.00 394.69 (8.3) 0.74 (0.0) 0.00 390.67 (8.2)
e5 0.80 (0.0) 0.53 13.38 (1.2) 0.03 (0.0) 0.00 400.85 (8.2) 0.75 (0.0) 0.00 399.78 (8.2)
gte gra. 0.73 (0.0) 0.94 1.33 (0.1) 0.00 (0.0) 0.00 408.81 (8.3) 0.56 (0.0) 0.01 398.16 (8.2)
gtr 0.71 (0.0) 0.95 1.29 (0.1) 0.04 (0.0) 0.00 386.58 (8.3) 0.71 (0.0) 0.00 383.41 (8.1)
ste. 0.86 (0.0) 1.00 1.00 (0.0) 0.58 (0.0) 1.00 1.00 (0.0) 1.00 (0.0) 1.00 1.00 (0.0)
e5 0.83 (0.0) 0.91 1.57 (0.2) 0.68 (0.0) 1.00 1.00 (0.0) 1.00 (0.0) 1.00 1.00 (0.0)
ste. gra. 0.79 (0.0) 0.99 1.09 (0.1) 0.00 (0.0) 0.00 418.16 (8.4) 0.57 (0.0) 0.00 399.56 (8.2)
gtr 0.77 (0.0) 1.00 1.00 (0.0) 0.00 (0.0) 0.00 393.07 (8.1) 0.71 (0.0) 0.00 390.26 (8.2)
gte 0.90 (0.0) 1.00 1.00 (0.0) 0.58 (0.0) 1.00 1.00 (0.0) 1.00 (0.0) 1.00 1.00 (0.0)
e5 0.85 (0.0) 0.98 1.05 (0.0) 0.37 (0.0) 0.89 1.55 (0.1) 1.00 (0.0) 1.00 1.00 (0.0)
e5 gra. 0.79 (0.0) 0.98 1.08 (0.0) 0.02 (0.0) 0.00 405.75 (8.3) 0.57 (0.0) 0.01 398.34 (8.2)
gtr 0.67 (0.0) 0.80 3.10 (0.6) 0.03 (0.0) 0.00 401.16 (8.4) 0.51 (0.0) 0.00 399.73 (8.2)
gte 0.87 (0.0) 0.99 1.02 (0.0) 0.68 (0.0) 1.00 1.00 (0.0) 1.00 (0.0) 1.00 1.00 (0.0)
ste. 0.75 (0.0) 0.98 1.06 (0.0) 0.37 (0.0) 1.00 1.00 (0.0) 1.00 (0.0) 1.00 1.00 (0.0)
@@
Table 6: Out-of-distribution translations on TweetTopic (with baselines): vec2vec models trained on NQ and evaluated on the entire TweetTopic test set (800 tweets). The rank metric varies from 1 to 800, thus 400 corresponds to a random ordering. Standard errors are shown in parentheses. Symbols denote the lowest-rank solver: Earth Mover’s Distance, Sinkhorn and Gromov-Wasserstein
vec2vec Naïve Baseline OT Baseline
$E_{1}$ $E_{2}$ $\cos(\cdot)$ T-1 Rank $\cos(\cdot)$ T-1 Rank $\cos(\cdot)$ T-1 Rank
gra. gtr 0.74 (0.0) 0.60 23.38 (1.6) -0.02 (0.0) 0.00 4010.00 (25.8) 0.82 (0.0) 0.00 3962.83 (26.1)
gte 0.85 (0.0) 0.08 346.21 (7.8) 0.01 (0.0) 0.00 3978.35 (26.1) 0.92 (0.0) 0.00 3808.18 (25.9)
ste. 0.72 (0.0) 0.13 242.23 (6.1) -0.01 (0.0) 0.00 3900.74 (26.2) 0.86 (0.0) 0.02 3780.44 (26.0)
e5 0.84 (0.0) 0.12 361.06 (8.7) 0.02 (0.0) 0.00 4024.92 (26.1) 0.93 (0.0) 0.00 3937.63 (26.2)
gtr gra. 0.78 (0.0) 0.51 35.27 (1.9) -0.02 (0.0) 0.00 4023.67 (26.1) 0.87 (0.0) 0.00 3964.83 (26.1)
gte 0.84 (0.0) 0.12 279.56 (6.9) 0.08 (0.0) 0.00 4180.47 (26.2) 0.87 (0.0) 0.00 4088.97 (26.2)
ste. 0.72 (0.0) 0.27 127.92 (4.4) 0.00 (0.0) 0.00 4296.04 (26.1) 0.76 (0.0) 0.00 4095.11 (26.1)
e5 0.82 (0.0) 0.01 1413.80 (18.3) 0.09 (0.0) 0.00 4064.47 (26.2) 0.93 (0.0) 0.00 4010.13 (26.1)
gte gra. 0.73 (0.0) 0.09 342.15 (7.8) 0.01 (0.0) 0.00 3946.19 (25.8) 0.87 (0.0) 0.00 3802.92 (25.9)
gtr 0.69 (0.0) 0.12 256.63 (6.4) 0.08 (0.0) 0.00 4229.90 (26.2) 0.69 (0.0) 0.00 4094.02 (26.1)
ste. 0.85 (0.0) 1.00 1.00 (0.0) 0.56 (0.0) 1.00 1.00 (0.0) 1.00 (0.0) 1.00 1.00 (0.0)
e5 0.86 (0.0) 0.54 17.71 (0.9) 0.69 (0.0) 0.98 1.04 (0.0) 1.00 (0.0) 1.00 1.00 (0.0)
ste. gra. 0.77 (0.0) 0.14 221.95 (5.9) -0.01 (0.0) 0.00 3951.42 (25.9) 0.87 (0.0) 0.01 3776.52 (26.0)
gtr 0.75 (0.0) 0.56 17.70 (1.0) 0.00 (0.0) 0.00 4339.83 (26.2) 0.70 (0.0) 0.00 4093.61 (26.1)
gte 0.91 (0.0) 1.00 1.00 (0.0) 0.56 (0.0) 1.00 1.00 (0.0) 1.00 (0.0) 1.00 1.00 (0.0)
e5 0.85 (0.0) 0.51 26.33 (1.2) 0.35 (0.0) 0.59 12.68 (0.6) 0.93 (0.0) 1.00 1.00 (0.0)
e5 gra. 0.78 (0.0) 0.21 151.09 (4.6) 0.02 (0.0) 0.00 4008.10 (25.9) 0.87 (0.0) 0.00 3932.58 (26.2)
gtr 0.66 (0.0) 0.01 1029.64 (14.9) 0.09 (0.0) 0.00 4032.85 (26.2) 0.82 (0.0) 0.00 4010.06 (26.1)
gte 0.87 (0.0) 0.60 32.59 (2.6) 0.69 (0.0) 0.98 1.09 (0.0) 1.00 (0.0) 1.00 1.00 (0.0)
ste. 0.75 (0.0) 0.46 32.12 (1.4) 0.35 (0.0) 0.86 2.49 (0.1) 0.86 (0.0) 1.00 1.01 (0.0)
@@
Table 7: Out-of-distribution translations on MIMIC (with baselines): vec2vec models trained on NQ and evaluated on an 8192-record subset of MIMIC. The rank metric varies from 1 to 8192, thus 4096 corresponds to a random ordering. Standard errors are shown in parentheses. Symbols denote the lowest-rank solver: Earth Mover’s Distance, Sinkhorn and Gromov-Wasserstein
Appendix D Ablations
D.1 Ablating components of our method
@@
Table 8: gte $\to$ gtr translators trained without individual components of our method on NQ and evaluated on a 65536-text subset of NQ (chunked in batches of 8192). The rank metric varies from 1 to 8192, thus 4096 corresponds to a random ordering. Standard errors are shown in parentheses.
D.2 Amount of data needed to learn translation
@@
Table 9: gte $\to$ gtr translators trained with different amounts of gte data: vec2vec models trained on NQ and evaluated an 8192-record subset of NQ. The rank metric varies from 1 to 8192, thus 4096 corresponds to a random ordering. Standard errors are shown in parentheses.
@@
Appendix E Prompt for measuring information extraction